# Variational Inference for Sparse Poisson Regression

### v9

@@ Acknowledgments
The authors would like to thank five anonymous reviewers for their valuable comments and corrections, which significantly improved the results of this paper. The reseach is based upon research funded by Iran National Science Foundation (INSF) under project No. 4048851.
@@
For low-dimensional settings, the recommended approach is consistent across objectives: the Bernoulli-VB or Laplace-VB methods are suggested for enforcing sparsity, performing parameter estimation, and generating predictions. In contrast, for high-dimensional problems, the recommendation shifts to favor the CS-VB and Laplace-VB methods.

### v10

@@ Acknowledgments
The authors would like to thank five anonymous reviewers for their valuable comments and corrections, which significantly improved the results of this paper.
@@
For low-dimensional settings, the recommended approach is consistent across objectives: the Bernoulli-VB or Laplace-VB methods are suggested for enforcing sparsity, performing parameter estimation, and generating predictions, because they have achieved the lowest median FPR, FNR, CRE, and TSRE in Figures 7 and 8. In contrast, for high-dimensional problems, since these criteria achieve their lowest medians for the CS-VB and Laplace-VB models, the recommendation shifts to favor these two methods (see also the additional results in the supplementary material).